# Structure growth in $f(Q)$ cosmology

### v1

@@ 4 The structure growth in $f(Q)$-cosmology
$f\equiv\frac{{\rm d}\ln{\mathcal{{D}}_{m}}}{{\rm d}\ln{a}}=-(1+z)\frac{1}{\mathcal{D}_{m}}\frac{{\rm d}\mathcal{D}_{m}(z)}{{\rm d}z}\;,$ (4.1)
Thus, by substituting the definition of (4.1) into the quasi-static second-order evolution equation (3.29), for the matter-dominated universe, the evolution of the growth rate is governed by the following expression [^9] For the case of $n=0$, Eq. (4.3) exactly reduced to $\Lambda$CDM. Indeed, by resorting to Eqs. (2.9) and (3.24) it is straightforward to obtain $(1+z)f^{\prime}=f^{2}-\left[(1+z)\frac{H^{\prime}}{H}-2\right]f-\frac{3\Omega_{m}}{2E^{2}}(1+z)^{3}\;.$
$(1+z)f^{\prime}$ $=$ $f^{2}-\left[(1+z)\frac{H^{\prime}}{H}+\left(1-\Omega_{m}\right)nE^{2n-2}-2\right]f$ (4.3)
-- $\,\frac{3\,\Omega_{m}(1+z)^{3}}{2\,\left[\Omega_{m}(1+z)^{3}+(1-\Omega_{m})E^{2n}\right]}\;.$
@@
$f\sigma_{8}(z)=-(1+z)\sigma_{8}(z)\frac{{\rm d}\mathcal{D}_{m}(z)}{{\mathcal{D}_{m}(z_{in})\rm d}z}\;,$ (4.4)
and $\sigma_{8}(z)$ for the given redshift $z_{in}$ can be expressed as [38]
$\sigma_{8}(z)=\sigma_{8}(z_{in})\frac{\mathcal{D}_{m}(z)}{\mathcal{D}_{m}(z_{in})}\;.$ (4.5)
In the following two consecutive subsections, we characterize and constrain the effect of our $f(Q)$-gravity model in the growth of structures, by admitting that the sets of observables $\sigma_{8}+f$, $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$ as explained in the Introduction - are all that we have available to compare with theoretical predictions issued from equations (4.3) - (4.5).
4.1 Constraining parameters
In this section, we take into account the growth structure data from Table 6 and 6 (see Appendix A), $\sigma_{8}+f$, $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$ to obtain the best-fit values of $\Omega_{m}$, $\sigma_{8}$, and $n$ through the MCMC simulations. These best-fit values of these parameters are also provided using the combinations of the cosmic expansion measurements with the growth structure data, meaning that OHD + SNIa +$f\sigma_{8}$, OHD + SNIa +$\sigma_{8}+f$, and OHD + SNIa +$\sigma_{8}+f+f\sigma_{8}$ datasets for both the $\Lambda$CDM model and the $f(Q)$ gravity power-law models. All constrained parameters are presented in Table 3. Furthermore, in Fig. 4 we provide the combined analysis of the constrained parameters using these three data sets $\sigma_{8}+f$, $f\sigma_{8}$, and $f\sigma_{8}+f+\sigma_{8}$, whereas in Fig. 5 we present the contour plots obtained when those growth data are used together with cosmic expansion measurements (i.e., OHD+SNIa+$f\sigma_{8}$, OHD+SNIa+$\sigma_{8}+f$ and OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$). In both figures 4 and 5, $1\sigma$ and $2\sigma$ levels are depicted and the corresponding $\Lambda$CDM contour plots appear on the left panels for comparison. A simple inspection of those two figures leads us to conclude that the consideration of the growth rate $f$ and the amplitude matter fluctuations $\sigma_{8}$ measurements from VIPERS and SDSS only (Fig. 4) or combined with OHD+SNIa (Fig. 5), significantly improves the constraints on the parameters of the model. Nonetheless, when standard redshift-space distortion data are considered too, such an improvement is reduced, although still present, particularly when OHD+SNIa are considered (Fig. 5, green and red contour remain of similar shape and area). Finally, by considering the best-fit parameters of $\Omega_{m}=0.265$, $n=0.020^{+0.005}_{-0.008}$ =$=\{0.025,0.02,0.012\}$ as obtained from the fitting against OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ data (see Table 3), we have plotted in Fig. 6 the evolution of density contrast $\delta(z)$ as governed by Eq. (3.29).
@@ 4.1 Constraining parameters
$f\sigma_{8}$ $0.253^{+0.020}_{-0.042}$ $0.804^{+0.035}_{-0.035}$
$\sigma_{8}+f$ $0.267^{+0.010}_{-0.010}$ $0.800^{+0.015}_{-0.015}$
$f+\sigma_{8}+f\sigma_{8}$ $0.258^{+0.025}_{-0.038}$ $0.803^{+0.036}_{-0.036}$
OHD+SNIa+$f\sigma_{8}$ $0.263^{+0.008}_{-0.008}$ $0.794^{+0.024}_{-0.024}$ $72.20^{+0.18}_{-0.18}$
OHD+SNIa+$\sigma_{8}+f$ $0.258^{+0.006}_{-0.007}$ $0.808^{+0.011}_{-0.011}$ $72.21^{+0.17}_{-0.17}$
OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ $0.262^{+0.010}_{-0.010}$ $0.796^{+0.024}_{-0.024}$ $72.14^{+0.19}_{-0.19}$
@@
$f\sigma_{8}$ $0.285^{+0.038}_{-0.053}$ $0.811^{+0.052}_{-0.052}$ $-0.069^{+0.089}_{-0.022}$
$\sigma_{8}+f$ $0.287^{+0.020}_{-0.027}$ $0.822^{+0.020}_{-0.025}$ $0.079^{+0.073}_{-0.073}$
$\sigma_{8}+f+f\sigma_{8}$ $0.280^{+0.031}_{-0.039}$ $0.804^{+0.039}_{-0.0146}$ $-0.001^{+0.038}_{-0.053}$
OHD+SNIa+$f\sigma_{8}$ $0.271^{+0.024}_{-0.021}$ $0.823^{+0.041}_{-0.033}$ $0.024^{+0.008}_{-0.010}$ $70.1^{+2.3}_{-3.3}$
OHD+SNIa+$\sigma_{8}+f$ $0.258^{+0.070}_{-0.070}$ $0.770^{+0.087}_{-0.087}$ $0.017^{+0.013}_{-0.005}$ $72.53^{+0.33}_{-0.30}$
OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ $0.265^{+0.075}_{-0.012}$ $0.802^{+0.024}_{-0.027}$ $0.020^{+0.005}_{-0.008}$ $71.50^{+1.20}_{-0.38}$
@@
Table 3: Best-fit values for cosmological parameters $\sigma_{8}$, $\Omega_{m}$, $H_{0}$ and $n$ for both $\Lambda$CDM and $f(Q)$-gravity power-law models using $f\sigma{{}_{8}}$, $\sigma_{8}+f$, $\sigma_{8}+f+f\sigma_{8}$, OHD+SNIa+$f\sigma_{8}$, OHD+SNIa+$\sigma_{8}+f$ and OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$. The value for the matter fluctuation amplitude ($\sigma_{8})$ is $\sigma_{8}=0.811\pm 0.006$ from Planck-2018 measurements of CMB [39].
Figure 6: Outer-Panel: Growth of density fluctuations for the matter-dominated universe as per Eq. (3.29), i.e., within the quasi-static approximation. Inner panel: redshift evolution for the deviation $\zeta(z)$ between $\Lambda$CDM and $f(Q)$-gravity model as per Eq. (3.31). For illustrative purposes, we use the best-fit values of $\Omega_{m}=0.265$, and $n=0.020^{+0.005}_{-0.008}=\{0.025,0.020,0.012\}$ obtained from OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ datasets, see Table 3.
@@ 4.2 Statistical analysis
Another crucial point in this section is to present the statistical values of ${\mathcal{L}(\hat{\theta}|data)}$, reduced $\chi^{2}$, $\chi^{2}$, $AIC$, $|\Delta AIC|$, $BIC$, and $|\Delta BIC|$ to establish the viability of the $f(Q)$ power-law model. As mentioned above, the $f(Q)$ gravity models will be tested against the the $\Lambda$CDM model, which will be considered as the benchmark accepted model. In order to do so, the growth structure data from Tables 6 and 6, as well as the cosmic expansion measurements from Section 2 have been considered for a comprehensive analysis. From the results as presented in Table 4, we show that the $f(Q)$ power-law model has observational support based on Jeffrey’s scale range for the analyses performed with $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$, OHD+SNIa+$f\sigma_{8}$, OHD+SNIa+$\sigma_{8}+f$, OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ datasets, whereas the model has less observational support for the cases of the $\sigma_{8}+f$ and OHD+SNIa+$\sigma_{8}+f$ datasets since the $|\Delta BIC|$ values fall within the range of less observational support.
In addition to constraining the best-fit values of the parameters, in Fig. 7 we have also presented the redshift-space distortion diagram $f\sigma_{8}$ using the evolution - quasi-static- equation Eq. (3.29) for matter-dominated universe. Therein, the redshift-space distortion best fittings are shown for both $\Lambda$CDM and $f(Q)$-gravity power-law models. To present the $f\sigma_{8}$ diagram, we consider the best-fit values $\sigma_{8}=0.804$, $\Omega_{m}=0.253$ for $\Lambda$CDM as given in Table 3. For the case of the $f(Q)$ gravity model, we have used different values of $n=-0.069^{+0.089}_{-0.022}=\{0.017,-0.069,-0.091\}$ with the corresponding values of $\sigma_{8}=0.811^{+0.038}_{-0.013}$ and $\Omega_{m}=0.285^{+0.038}_{-0.053}$. For the two latter parameters - $\sigma_{8}$ and $\Omega_{m}$- we have taken the central values. All these values are taken from the results presented in Table 3.
data $\mathcal{L}(\hat{\theta}|)$ $\chi^{2}$ $\chi^{2}$-red $AIC$ $|\Delta AIC|$ $BIC$ $|\Delta BIC|$
$\Lambda$CDM
$f\sigma_{8}$8.0 16 0.57 20 0 24.8 0
$\sigma_{8}+f$1.2 2.4 0.6 6.4 0 5.98 0
$\sigma_{8}+f+f\sigma_{8}$9.1 18.2 0.54 22.2 0 25.6 0
OHD+SNIa+$f\sigma_{8}$540.8 1081.6 0.98 1085.6 0 1095.63 0
OHD+SNIa+$\sigma_{8}+f$532.2 1064.4 0.98 1068.4 0 1078.4 0
OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$542.1 1084.2 0.986 1088.2 0 1098.30 0
$f(Q)$ model
$f\sigma_{8}$8.6 17.20 0.65 23.20 3.20 27.40 3.30
$\sigma_{8}+f$2.40 4.80 0.93 10.80 4.40 10.17 4.20
$\sigma_{8}+f+f\sigma_{8}$9.00 18 0.55 24.00 1.80 28.99 3.38
OHD+SNIa+$f\sigma_{8}$538.10 1076.20 0.96 1082.40 2.20 1097.26 1.63
OHD+SNIa+$\sigma_{8}+f$531.50 1063.00 0.98 1069.00 2.60 1084.06 5.57
OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$539.50 1080.0 0.97 1086.0 2.0 1101.50 3.20
Table 4: Values of ${\mathcal{L}(\hat{\theta}|data)}$, $\chi^{2}$, $\chi^{2}$-red(uced), $AIC$, $|\Delta AIC|$, $BIC$ and $|\Delta BIC|$ for $\Lambda$CDM and $f(Q)$-gravity models. The $|\Delta AIC|$ values are lower than $3.6$ for all data sets whereas the $|\Delta BIC|$ values have been found to lie within the range of observational support according to Jeffrey’s scale, except for the cases of $\sigma_{8}+f$ and OHD+SNIa+$\sigma_{8}+f$.
Figure 7: Linear growth-structure $f\sigma_{8}$ diagrams for $\Lambda$CDM and $f(Q)$-gravity power-law models - for different values of $n$ with $f\sigma_{8}$ data. For illustrative purposes, we use $n=-0.069^{+0.089}_{-0.022}=\{0.02,-0.069,-0.091\}$ with the corresponding values of $\sigma_{8}=0.811^{+0.038}_{-0.013}$ and $\Omega_{m}=0.285^{+0.038}_{-0.053}$ as shown in Table 3. For the latter two parameters, $\sigma_{8}$ and $\Omega_{m}$- we have taken the central values.
5 Results and discussion

### v4

@@ 4 Structure growth in $f(Q)$ cosmology
$f\equiv\frac{{\rm d}\ln{\mathcal{{D}}_{m}}}{{\rm d}\ln{a}}=-(1+z)\frac{1}{\mathcal{D}_{m}}\frac{{\rm d}\mathcal{D}_{m}(z)}{{\rm d}z}\;.$ (48)
Thus, by substituting the definition of (48) into the quasi-static second-order evolution equation (3.1), for the matter-dominated universe, the evolution of the growth rate is governed by the following expression [^7] For the case of $n=0$, Eq. (4) exactly reduced to $\Lambda$CDM. Indeed, by resorting to Eqs. (9) and (39) it is straightforward to obtain $(1+z)f^{\prime}=f^{2}-\left[(1+z)\frac{H^{\prime}}{H}-2\right]f-\frac{3\Omega_{m}}{2E^{2}}(1+z)^{3}\;.$
$(1+z)f^{\prime}=f^{2}-\left[(1+z)\frac{H^{\prime}}{H}+\left(1-\Omega_{m}\right)nE^{2n-2}-2\right]f$
$-\,\frac{3\,\Omega_{m}(1+z)^{3}}{2\,\left[\Omega_{m}(1+z)^{3}+(1-\Omega_{m})E^{2n}\right]}\;.$ (50)
@@
$f\sigma_{8}(z)=-(1+z)\sigma_{8}\frac{\mathcal{D}^{\prime}_{m}(z)}{\mathcal{D}_{m}(z)}\;.$ (51)
In the following two consecutive subsections, we characterize and constrain the effect of our $f(Q)$-gravity model in the growth of structures, by admitting that the sets of observables $\sigma_{8}+f$, $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$ as explained in the Introduction - are all that we have available to compare with theoretical predictions issued from equations (4) and (51).
Figure 6: Outer-Panel: Growth of density fluctuations for the matter-dominated universe as per Eq. (3.1), i.e., within the quasi-static approximation. Inner panel: redshift evolution for the deviation $\zeta(z)$ between $\Lambda$CDM and $f(Q)$-gravity model as per Eq. (46). For illustrative purposes, we use the best-fit values of $\Omega_{m}=0.327$, and $n=-0.025^{+0.021}_{-0.010}=\{-0.004,-0.025,-0.035\}$ obtained from OHD + SNIa + $\sigma_{8}+f\,+\,f\sigma_{8}$ datasets, see Table 3.
Figure 7: Linear growth-structure $f\sigma_{8}$ diagrams for $\Lambda$CDM and $f(Q)$-gravity power-law models - for different values of $n$ with $f\sigma_{8}$ data. For illustrative purposes, we use $n=0.090^{+0.035}_{-0.012}=\{0.125,0.09,0.078\}$ with the corresponding values of $\sigma_{8}=0.792^{+0.012}_{-0.012}$ and $\Omega_{m}=0.300^{+0.037}_{-0.022}$ as shown in Table 3. For the latter two parameters, $\sigma_{8}$ and $\Omega_{m}$- we have taken the central values.
Figure 8: Comparison of $\Lambda$CDM and $f(Q)$-gravity model $H_{0}$ values (in km/s/Mpc) with various measurements.
[…]
@@
$f\sigma_{8}$ $0.288^{+0.023}_{-0.022}$ $0.806^{+0.019}_{-0.018}$
$\sigma_{8}+f$ $0.278^{+0.004}_{0.004}$ $0.780^{+0.009}_{0.009}$
$\sigma_{8}+f+f\sigma_{8}$ $0.290^{+0.009}_{0.009}$ $0.790^{+0.013}_{-0.013}$
OHD + SNIa + $f\sigma_{8}$ $0.323^{+0.009}_{-0.009}$ $0.800^{+0.015}_{-0.015}$ $73.09^{+0.60}_{-0.60}$
OHD + SNIa + $\sigma_{8}+f$ $0.321^{+0.007}_{-0.007}$ $0.815^{+0.011}_{-0.011}$ $73.43^{+0.82}_{-0.75}$
OHD + SNIa + $\sigma_{8}+f+f\sigma_{8}$ $0.336^{+0.007}_{-0.007}$ $0.801^{+0.008}_{-0.008}$ $72.94^{+0.17}_{-0.19}$
@@
$f\sigma_{8}$ $0.300^{+0.037}_{-0.022}$ $0.792^{+0.012}_{-0.012}$ $0.090^{+0.035}_{-0.012}$
$\sigma_{8}+f$ $0.275^{+0.020}_{-0.065}$ $0.774^{+0.009}_{0.009}$ $-0.120^{+0.013}_{-0.011}$
$\sigma_{8}+f+f\sigma_{8}$ $0.282^{+0.033}_{-0.021}$ $0.785^{+0.012}_{-0.012}$ $0.100^{+0.034}_{-0.011}$
OHD + SNIa + $f\sigma_{8}$ $0.335^{+0.058}_{-0.051}$ $0.809^{+0.012}_{-0.019}$ $-0.01^{+0.019}_{-0.018}$ $72.75^{+2.20}_{-2.50}$
OHD + SNIa + $\sigma_{8}+f$ $0.320^{+0.047}_{-0.039}$ $0.821^{+0.034}_{-0.034}$ $-0.013^{+0.010}_{-0.009}$ $73.32^{+2.70}_{-3.90}$
OHD + SNIa + $\sigma_{8}+f+f\sigma_{8}$ $0.327^{+0.056}_{-0.045}$ $0.826^{+0.026}_{-0.037}$ $-0.025^{+0.021}_{-0.010}$ $72.41^{+2.40}_{-3.10}$
@@
Table 3: Best-fit values for cosmological parameters $\sigma_{8}$, $\Omega_{m}$, $H_{0}$ and $n$ for both $\Lambda$CDM and $f(Q)$-gravity power-law models using $f\sigma{{}_{8}}$, $\sigma_{8}+f$, $\sigma_{8}+f\,+\,f\sigma_{8}$, OHD + SNIa + $f\sigma_{8}$, OHD + SNIa + $\sigma_{8}+f$ and OHD + SNIa + $\sigma_{8}+f\,+\,f\sigma_{8}$. The CMB Planck-2018 value for the matter fluctuation amplitude ($\sigma_{8})$ is $\sigma_{8}=0.811\pm 0.006$ (Aghanim et al., 2020).
4.1 Constraining parameters
In this section, we take into account the growth structure data from Table 6 and 6 (see Appendix A), $\sigma_{8}+f$, $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$ to obtain the best-fit values of $\Omega_{m}$, $\sigma_{8}$, and $n$ through the MCMC simulations. These best-fit values of these parameters are also provided using the combinations of cosmic expansion measurements with the growth structure data, meaning that the OHD + SNIa +$f\sigma_{8}$, OHD + SNIa +$\sigma_{8}+f$, and OHD + SNIa +$\sigma_{8}+f+f\sigma_{8}$ datasets for both the $\Lambda$CDM model and the $f(Q)$ gravity power-law models. All constrained parameters are presented in Table 3. Furthermore, in Fig. 4 we provide the combined analysis of the constrained parameters using these three datasets $\sigma_{8}+f$, $f\sigma_{8}$, and $f\sigma_{8}+f+\sigma_{8}$, whereas in Fig. 5 we present the contour plots obtained when those growth data are used together with cosmic expansion measurements (i.e., OHD + SNIa + $f\sigma_{8}$, OHD + SNIa + $\sigma_{8}+f$ and OHD + SNIa + $\sigma_{8}+f\,+\,f\sigma_{8}$). In both figures 4 and 5, $1\sigma$ and $2\sigma$ levels are depicted, and the corresponding $\Lambda$CDM contour plots appear on the left panels for comparison. A simple inspection of those two figures leads us to conclude that the consideration of the growth rate $f$ and the amplitude matter fluctuations $\sigma_{8}$ measurements from VIPERS and SDSS only (Fig. 4) or combined with OHD + SNIa (Fig. 5), significantly improves the constraints on the parameters of the model. Nonetheless, when standard redshift-space distortion data are considered too, such an improvement is reduced, although still present, particularly when OHD + SNIa are considered (Fig. 5, green and blue contours remain of similar shape and area). Finally, by considering the best-fit parameters of $\Omega_{m}=0.327$, and $n=-0.025^{+0.021}_{-0.010}=\{-0.004,-0.025,-0.035\}$ as obtained from the fitting against OHD + SNIa + $\sigma_{8}+f\,+\,f\sigma_{8}$ data (see Table 3), we have plotted in Fig. 6 the evolution of density contrast $\delta(z)$ as governed by Eq. (3.1).
In addition to constraining the best-fit values of the parameters, in Fig. 7 we have also presented the redshift-space distortion diagram $f\sigma_{8}$ using the evolution - quasi-static- equation Eq. (3.1) for the matter-dominated universe. Therein, the redshift-space distortion best fittings are shown for both $\Lambda$CDM and $f(Q)$-gravity power-law models. To present the $f\sigma_{8}$ diagram, we consider the best-fit values $\sigma_{8}=0.806$, $\Omega_{m}=0.288$ for $\Lambda$CDM as given in Table 3. For the case of the $f(Q)$ gravity model, we have used different values of $n=0.090^{+0.035}_{-0.012}=\{0.125,0.090,0.078\}$ with the corresponding values of $\sigma_{8}=0.792^{+0.012}_{-0.012}$ and $\Omega_{m}=0.300^{+0.037}_{-0.022}$. For the two latter parameters - $\sigma_{8}$ and $\Omega_{m}$ we have taken the central values. All these values are taken from the results presented in Table 3. For both $\Lambda$CDM and $f(Q)$-gravity models, in Fig. 9, we compare the $H_{0}$ values as obtained from the joint analyses as shown in Table 3 with different current measurements. The comparisons are made resorting to $i)$ early measurements, including Planck 2018 (Aghanim et al., 2020), Dark Energy Spectroscopy Instrument (DESI) (Adame et al., 2024); and $ii)$ late-time measurements derived from direct observations of the local Universe, such as Supernovae and $H_{0}$ for the Equation of State (SH0ES) (Riess et al., 2019), H0LiCOW (Wong et al., 2020), and Hubble Space Telescope (HST) (Riess et al., 2011). Figure 9 shows the discrepancy of $H_{0}$ measurements from early-universe evidence, late-universe observations, and the theoretical models $\Lambda$CDM and $f(Q)$ studied herein. The usual $\Lambda$CDM model corresponds well with the late measurements, which predict a higher $H_{0}$, while late-time measurements from sources like SH0ES and JWST consistently provide larger values. The $f(Q)$ model yields higher $H_{0}$ values, suggesting a better agreement with late observations, while it does not fully resolve the discrepancy of $H_{0}$ measurements.
Analogously, the comparison of $\sigma_{8}$ results with different experimental findings is showcased in Fig. 9 resorting to $i)$ low-redshift measurements ($z\leq 1$) namely HSC (Hamana et al., 2020), and KiDS-450 (Hildebrandt et al., 2017); and $ii)$ high-redshift measurements ($z>1$) namely DESI (Adame et al., 2024), and Planck 2018 (Aghanim et al., 2020). Such a figure shows that, with the exception of the $\sigma_{8}+f$ datasets, the best-fit values of $\sigma_{8}$ found with our analyses remain close to the high-redshift measurements for both $\Lambda$CDM and $f(Q)$ gravity models.
Data $\mathcal{L}(\hat{\theta}|{\rm data})$ $\chi^{2}$ $\chi^{2}$-red $\rm{AIC}$ $\Delta\rm{AIC}$ $\rm{BIC}$ $\Delta\rm{BIC}$
$\Lambda$CDM
$f\sigma_{8}$8.00 16.00 0.57 20 24.80
$\sigma_{8}+f$1.20 2.40 0.60 6.40 5.98
$\sigma_{8}+f+f\sigma_{8}$9.10 18.20 0.54 22.20 25.60
OHD+SNIa+$f\sigma_{8}$791.24 1582.48 0.93 1588.48 1605.93
OHD+SNIa+$\sigma_{8}+f$783.83 1567.66 0.87 1573.66 1590.04
OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$793.89 1587.78 0.91 1593.78 1610.23
$f(Q)$ model
$f\sigma_{8}$8.90 17.80 0.75 23.80 3.80 28.00 3.20
$\sigma_{8}+f$1.82 3.64 0.91 9.64 3.24 7.00 1.21
$\sigma_{8}+f+f\sigma_{8}$9.04 18.08 0.60 24.08 1.88 28.46 3.04
OHD+SNIa+$f\sigma_{8}$790.15 1581.04 0.92 1589.04 0.56 1610.96 5.03
OHD+SNIa+$\sigma_{8}+f$783.05 1566.10 0.86 1574.1 0.45 1595.96 5.92
OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$793.450 1586.90 0.89 1594.90 1.12 1616.06 5.82
Table 4: Values of ${\mathcal{L}(\hat{\theta}|data)}$, $\chi^{2}$, $\chi^{2}$-red, $\rm{AIC}$, $\Delta\rm{AIC}$, $\rm{BIC}$ and $\Delta\rm{BIC}$ for $\Lambda$CDM and $f(Q)$-gravity models. The $\Delta\rm{AIC}$ values are lower than $3.80$ for all datasets, whereas the $\Delta\rm{BIC}$ values have been found to lie within the range of observational support, except for the case of OHD+SNIa+$\sigma_{8}+f$, the latter indicating less observational support. Interestingly, the OHD+SNIa+$f\sigma_{8}$ analysis indicates strong support when resorting to both criteria. Then, the addition of just three $\sigma_{8}+f$ data to OHD+SNIa+$f\sigma_{8}$ reduces the support, as per $\Delta$BIC, significantly although according to $\Delta$AIC, the dataset OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ still remains strongly supported.
@@ 4.2 Statistical analysis
Another crucial point in this section is to present the statistical values of ${\mathcal{L}(\hat{\theta}|{\rm data})}$, reduced $\chi^{2}$, $\chi^{2}$, $\rm{AIC}$, $\Delta\rm{AIC}$, $\rm{BIC}$, and $\Delta\rm{BIC}$ to establish the viability of the $f(Q)$ power-law model. As mentioned above, the $f(Q)$ gravity models will be tested against the $\Lambda$CDM model, which will be considered as the benchmark accepted model. In order to do so, the growth structure data from Tables 6 and 6, as well as the cosmic expansion measurements from Section 2 have been considered for a comprehensive analysis. From the results presented in Table 4 we conclude that the $f(Q)$ power-law model has observational support based on our model selection criteria. In particular, $\Delta$AIC values lie between 0.34 and 3.80 for $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$, OHD+SNIa+$f\sigma_{8}$, OHD+SNIa+$\sigma_{8}+f$, OHD+SNIa+$\sigma_{8}+f+f\sigma_{8}$ datasets. Also, $\Delta$BIC values are less than 6, this indicates that the $f(Q)$ gravity model has ranges positively against $\Lambda$CDM, i.e., not completely ruled out.
5 Conclusions
In the context of the nonmetric modified theory of gravity, we have performed a thorough statistical analysis resorting to both cosmological expansion history and large-scale structure data. Throughout the investigation, we considered a generic class of $f(Q)$ power-law models whose gravitational Lagrangian is of the form $f(Q)=Q+\alpha\,Q^{n}$.